### v2

# Extracting the Speed of Sound in Heavy-Ion Collisions: A Study of Quantum Fluctuations and Thermalization

Abstract
The thermalization of quark-gluon plasma created in heavy-ion collisions is crucial for understanding its behavior as a relativistic fluid and the thermodynamic properties of the Quantum Chromodynamics (QCD). This study investigates the role of fluctuations in the relationship between transverse momentum and particle multiplicity, with a particular focus on their impact on extracting the QCD speed of sound. In a thermalized quark-gluon plasma, fluctuations are dominantly quantum in origin and follow a Gaussian distribution due to their independence from the thermodynamic response. In contrast, non-thermalized systems display non-Gaussian fluctuations, reflecting the breakdown of thermalization. By leveraging the Gaussianity condition of quantum fluctuations, the physical value of the speed of sound can be extracted statistically, even in the presence of significant event-by-event fluctuations. This framework provides a robust diagnostic tool for probing thermalization and extracting thermodynamic properties in both large and small collision systems.
Heavy-ion collisions at facilities such as the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) create extreme conditions, where Quark-Glasma Plasma (QGP) forms Shuryak (2017). A crucial question is whether this deconfined state of quarks and gluons achieves local thermal equilibrium, at least transiently, so that one is allowed to apply hydrodynamic models Gale et al. (2013); Romatschke and Romatschke (2019); Florkowski et al. (2018) and explore the phase structure of the strongly interacting matter through fundamental thermodynamic properties Luo and Xu (2017). As a closed quantum system of high energies, the thermalization of QGP is also essential for understanding non-equilibrium dynamics, in a broad sense Schlichting and Teaney (2019); Shen and Yan (2020); Berges et al. (2021); Brauner et al. (2022).
Local thermalization of QGP in nucleus-nucleus collisions is widely accepted, supported by different types of flow signatures that reveal anisotropic particle emissions. These are indirect evidence, however, as they are consequences of hydrodynamic response of a fluid-like QGP with respect to the geometric deformation of initial state Heinz and Snellings (2013); Yan (2018). More explicitly, by assuming the QGP as a relativistic fluid, the thermodynamic quantities like pressure $P$, energy density $e$, and entropy $S$ can be well implemented in hydrodynamic models, and the initial geometric deformation results in anisotropies of pressure gradients, which further drives the system to expand anisotropically Ollitrault (1992). Although flow analyses have extended to the smaller colliding systems, including proton-lead Abelev et al. (2013); Aad et al. (2013); Khachatryan et al. (2015), proton-proton Khachatryan et al. (2010); Aad et al. (2016a); Acharya et al. (2024a) and even electron-positron Chen et al. (2024), aside from which, investigation of the local thermalization of the strongly interacting matter through directly the thermodynamic properties, such as the QCD speed of sound Van Hove (1982), was realized only recently Gardim et al. (2020a, b).
Among the thermodynamic properties of QCD, the speed of sound, $c_{s}$, holds special significance as it contains information of the equation of state (EOS). Defined as
$c_{s}^{2}\equiv\frac{\partial P}{\partial e}=\frac{\partial\ln T}{\partial\ln S}\,,$ (1)
the speed of sound measures the thermodynamic response of pressure to energy density, as well as the response of relative variation of temperature to the relative variation of entropy. Note that the second equation applies to a uniform and baryon-less QGP. Speed of sound is closely linked to the stiffness of the EOS. Near the critical region, $c_{s}^{2}$ is expected to exhibit a sharp decrease, signaling a softening of the EOS due to the transition between hadronic matter and QGP Bazavov et al. (2014).
In high-energy ultracentral collisions of heavy nuclei, where geometric deformation is substantially suppressed, if the created QGP reaches local thermal equilibrium, the QCD speed of sound governs the response of variation in the observed mean transverse momentum of particles, $\langle p_{T}\rangle\equiv\sum p_{T}/N_{\rm c}$, to the variation in the charged particle multiplicity, $N_{\rm c}$. This is because, in ultracentral collisions, mean transverse momentum measures effectively temperature in the system Gardim et al. (2020a), while the charged particle multiplicity is linked to the entropy. Consequently, linear relationships exist between $\langle p_{T}\rangle$ and $T$, and between $N_{\rm c}$ and $S$. Corresponding to Eq. (1), the response is purely thermodynamic and deterministic, so one has Gardim et al. (2020a, b)
$\frac{\Delta_{p}}{\langle p_{T}\rangle_{0}}=c_{s}^{2}\frac{\Delta_{N}}{N_{0}}\,,$ (2)
where $\langle p_{T}\rangle_{0}$ and $N_{0}$ are the averaged values of the mean transverse momentum and the charged particle multiplicity over the ultracentral collision events, while $\Delta_{p}=\langle p_{T}\rangle-\langle p_{T}\rangle_{0}$ and $\Delta_{N}=N_{\rm c}-N_{0}$ are variations. Note that the mechanism of relating $\Delta_{p}$ and $\Delta_{N}$ is not unique. From purely multi-parton interactions or mini-jet production, for instance, a linear response relation between $\Delta_{p}$ and $\Delta_{N}$ presents analogously, but the coefficient has of course, nothing to do with the speed of sound. Nonetheless, it is crucial that Eq. (2) from a thermalized QGP is thermodynamic, while the correlations induced in non-thermal systems are quantum.
Figure 1: The normalized two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{N},\Delta_{p})$, for the central Pb-Pb (0-5%) and p-Pb (0-2%) collisions at $\sqrt{s_{NN}}=5.02$ TeV, from simulations using hydrodynamics, HIJING and PYTHIA. The growth of width of the distributions, from (a) to (d), indicating the increasing of fluctuation strength. The tilted tip of the distributions (except PYTHIA) implies a positive correlations between $\Delta_{p}$ and $\Delta_{N}$.
While in experiments, a linear rise between $\Delta_{p}$ and $\Delta_{N}$ has been identified in the ultracentral lead-lead collisions at $\sqrt{s_{NN}}$ = 5.02 TeV Hayrapetyan et al. (2024); Acharya et al. (2024b), the extracted speed of sound from the slope is contaminated. In several model studies using hydrodynamic simulations, it has been noticed that the slope depends on the selection of rapidity, cut in transverse momentum, and how centrality is determined Soares Rocha et al. (2024); Nijs and van der Schee (2024). These kinematic effects can be understood, at least qualitatively. For instance, recording in a finite interval of the transverse momentum of particle spectrum reduces the number of thermal particles taken into account, which as a consequence changes the contribution of thermodynamic response.
On the other hand, there are fluctuations in realistic heavy-ion collisions. On an event-by-event basis, these fluctuations are dominantly of quantum origin, such as the quantum fluctuations of the nucleons populating inside nuclei as they collide, and decay and scatterings of hadrons during final-state particlization the . In terms of flow measurement, fluctuations in high-energy heavy-ion collisions have been well captured through multi-particle correlations Shen and Yan (2020), in consistent to hydrodynamic model simulations on an event-by-event basis. The role of fluctuations in the response relation between $\Delta_{p}$ and $\Delta_{N}$, in the ultracentral collisions, remains unclear so far Sun and Yan (2024). In this Letter, we will investigate the effect of fluctuations on the extraction of speed of sound, and we propose a statistical method that systematically isolates the thermodynamic response, so that in both theoretical simulations and experiments the speed of sound can be well identified in the presence of the fluctuations. More importantly, we find that the quantum nature of the event-by-event fluctuations makes the extraction of the speed of sound an ideal probe of the QGP thermalization.
Quantum fluctuations. – In realistic heavy-ion collisions, with respect to the mean transverse momentum and the charged particle multiplicity, the presence of fluctuations leads to a two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{p},\Delta_{N})$ Samanta et al. (2024, 2023); Aad et al. (2024), instead of the deterministic relation Eq. (2).
While such a distribution can be measured experimentally, for the purpose of illustration, we generate a set of $\mathcal{P}(\Delta_{p},\Delta_{N})$ from numerical simulations of hydrodynamic model and non-thermal models num . The hydrodynamic model is hybrid that incorporates the generation of initial density profile using the TRENTo method Moreland et al. (2015), solving coupled equations of viscous hydrodynamics with the EOS from lattice QCD (LEOS) Bazavov et al. (2014) for the system’s dynamical evolution using the MUSIC program Schenke et al. (2011, 2010); Paquet et al. (2016), which is then connected to UrQMD for particlization and hadron gas evolution Bass et al. (1998). These are standard theoretical prescriptions that systematically takes into account the event-by-event quantumn fluctuations from different stages. It should be emphasized that while solving hydrodynamics implies local thermalization up to the dissipative corrections, the hadron dynamics regarding UrQMD does not require local thermalization. Therefore, local thermalization from the numerical simulations of the hybrid hydrodynamic models is only transiently realized.
For lead-lead (Pb-Pb) collisions at high energies, longitudinal boost invariance is a good approximation, which allows us to simplify the analysis and work in a 2+1 dimensional setup. However, this longitudinal symmetry is broken in small collision systems such as proton-lead (p-Pb), for which we invoke the exact 3+1 dimensional characterization of the initial state in TRENTo-3D Soeder et al. (2023), as well as in MUSIC. For comparison, we also simulate and collect data on the mean transverse momentum and multiplicity in p-Pb collisions using non-thermal models, including HIJING Wang and Gyulassy (1991); Deng et al. (2011); Wang and Gyulassy (1992) and PYTHIA Sjostrand et al. (2008); Sjöstrand et al. (2015); Bierlich et al. (2022). In these non-thermal models, particles are generated by parton scatterings and radiative processes rather than thermal medium expansion, meaning that local thermalization is absent.
Figure 2: The probability distribution $\rho(\delta|c_{s}^{2})$ with $c_{s}^{2}$ determined by the zero skewness condition $\{\delta^{3}\}_{c}=0$ (red histogram) and $\{\delta^{5}\}_{c}=0$ (blue histogram). For comparison, the Gaussian distributions with zero mean and the variance $\{\delta^{2}\}^{1/2}$ determined with respect to the extracted $c_{s}^{2}$ values are shown by lines of the same colors.
With the $\langle p_{T}\rangle_{0}$ and $N_{0}$ fixed with respect to the top central collisions in experiments, for instance, top 5% centrality for Pb-Pb collisions at the LHC with respect to the pseudo-rapidity $|\eta|<0.5$ and particles of transverse momentum $p_{T}\in[0,3]$ GeV Aad et al. (2016b); Acharya et al. (2023), from numerical simulations we obtain the two-dimensional joint probability distributions, as presented in Fig. 1. In Pb-Pb collisions, the vanishing impact parameter, or saturation of system size, appears only in the ultracentral events, namely, the right tip of the distribution in Fig. 1 (a) Samanta et al. (2023); Gardim et al. (2020b). In p-Pb collisions, on the other hand, the geometry of the generated system is dominantly determined by the proton, one does not require the condition of ultracentral collisions to suppress the size dependence Gardim et al. (2024). In principle, the analysis of the speed of sound can be carried out through most centralities of the p-Pb collisions. Nonetheless, we still focus on the top 2% central events in p-Pb, where system collective expansion has been observed in experiments through particle flow of multi-particle correlations Khachatryan et al. (2015).
For later convenience, we have also obtained the effective temperatures from the freeze-out hypersurface Gardim et al. (2020a), with details given in the Supplemental Materials. While for the Pb-Pb system, one approximately has $\langle p_{T}\rangle\approx 3T_{\rm eff}$ due to the boost invariance, for p-Pb collisions, on an event-by-event basis we find that $\langle p_{T}\rangle\approx 2.22\sim 2.8\,T_{\rm eff}$ and $\langle p_{T}\rangle\approx 2.16\sim 2.58\,T_{\rm eff}$ for the collision energies $\sqrt{s_{NN}}=5.02$ TeV and $8.16$ TeV, respectively.
Given the joint probability distribution $\mathcal{P}(\Delta_{p},\Delta_{N})$, the response between the variation of the transverse momentum and the variation of multiplicity should be generalized with a stochastic contribution, $\delta$,
$\frac{\Delta_{p}}{\langle p_{T}\rangle_{0}}=c_{s}^{2}\frac{\Delta_{N}+\delta}{N_{0}}\,.$ (3)
Importantly, for a thermalized QGP, $\delta$ is independent of the thermodynamic response relation. This independence reflects the quantum origin of $\delta$, which introduces stochastic fluctuations uncorrelated with the deterministic macroscopic relationship. While $\delta$ is independent of the thermodynamic response, it may exhibit statistical correlations with individual observables like $\Delta_{N}$ or $\Delta_{p}$, arising from shared quantum or initial-state fluctuations. In a thermalized system, where $\delta$ varies randomly and independently across events, its distribution approaches Gaussianity by the Central Limit Theorem (CLT) Samanta et al. (2024, 2023). On the other hand, in a system without thermalization, the response relation itself (where the response coefficient is no longer the speed of sound) is quantum, hence $\delta$ is not independent from event to event and the $\delta$-distribution is generically non-Gaussian. This non-Gaussinity in $\delta$ in non-thermal systems can be conceptually understood as a consenquence of interactions among these fluctuations, through the quantum response relation. This is analogous to the non-Gaussianity observed in primordial quantum fluctuations, which signals interactions in the early universe Maldacena (2003).
Irrespective of the condition of local thermalization, finite-size effects, proportional to $1/N_{0}$, introduce residual non-Gaussianity to the $\delta$-distribution. Specifically, corrections to higher-order moments of the $\delta$-distribution scale as powers of $1/N_{0}$, with the skewness scaling as $1/N_{0}^{3/2}$, and kurtosis scaling as $1/N_{0}^{2}$, etc. Thus, the previous discussions of Gaussianity in thermlized QGP and non-Gaussianity in non-thermal systems apply up to corrections of order $1/N_{0}^{a}$, where $a$ depends on the specific observable or moment. These finite-size effects are particularly relevant when $N_{0}$ is relatively small.
To verify the analysis, we need to solve the distribution of $\delta$ with respect to the physical value of the speed of sound. Regarding Eq. (3), it is a mathematical problem: How to solve the constant parameter $c_{s}^{2}$, given two known distributions of $\Delta_{p}$ and $\Delta_{N}$, and one unknown distribution of $\delta$? In fact, Eq. (3) implicitly defines the conditional distribution, $\rho(\delta|c_{s}^{2})$. In the small $c_{s}^{2}$ limit, $\rho(\delta|c_{s}^{2})$ approaches the distribution of $\Delta_{p}/\langle p_{T}\rangle_{0}$, while for $c_{s}^{2}\to\infty$, the distribution of $\delta$ is equivalent to that of $-\Delta_{N}/N_{0}$.
[…]
$(c_{s}^{2})^{3}\frac{\{\Delta_{N}^{3}\}}{N_{0}^{3}}-3(c_{s}^{2})^{2}\frac{\{\Delta_{N}^{2}\Delta_{p}\}}{N_{0}^{2}\langle p_{T}\rangle_{0}}+3c_{s}^{2}\frac{\{\Delta_{N}\Delta_{p}^{2}\}}{N_{0}\langle p_{T}\rangle_{0}^{2}}-\frac{\{\Delta_{p}^{3}\}}{\langle p_{T}\rangle_{0}^{3}}=0\,,$ (4)
with the physical value of the speed of sound its real root. The coefficients of the equation correspond respectively, to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{p},\Delta_{N})$, which are all measurable in realistic experiments. With the speed of sound extracted from the zero skewness condition, one is allowed to solve the exact probability distribution $\rho(\delta|c_{s}^{2})$. As shown in Fig. 2, to a good approximation within corrections of $1/N_{0}^{a}$, we find that $\delta$ from hydrodynamic simulations follow the Gaussian distribution, while from non-thermal simulations, fluctuations of $\delta$ exhibit apparently non-Gaussianity with heavy tails. We have also solved similarly, with respect to the condition that the fifth order cumulant $\{\delta^{5}\}_{c}=\{\delta^{5}\}-10\{\delta^{2}\}\{\delta^{3}\}=0$, which gives consistent results. Besides, for the values of $c_{s}^{2}$ other than the desired ones, we notice that even from hydrodynamic models $\rho(\delta|c_{s}^{2})$ becomes non-Gaussian.
Extracting speed of sound in the presence of fluctuations. – Without any concern of the effect of fluctuations, Eq. (3) can be solved via a sub-bin selection of the ultracentral events, as has been carried out so far in experiments Hayrapetyan et al. (2024); Acharya et al. (2024b) and theoretical model studies Soares Rocha et al. (2024); Nijs and van der Schee (2024). After event average in the sub-bins, the variation of the mean transverse momentum and multiplicity can be quantified between sub-bins, and one has
$\frac{\{\Delta_{p}\}_{I}}{\langle p_{T}\rangle_{0}}=c_{s}^{2}\frac{\{\Delta_{N}\}_{I}+\{\delta\}_{I}}{N_{0}}\,,\quad I=1,2,\ldots$ (5)
where {}$\{\ldots\}_{I}$ denotes event average in the sub-bin $I$. If $\delta$ and $\Delta_{N}$ are independent in the sub-bins, the speed of sound can be simply read off as the slope from the resulted line between $\{\Delta_{p}\}_{I}/\langle p_{T}\rangle_{0}$ and $\{\Delta_{N}\}_{I}/N_{0}$, otherwise, the slope gets correction from $\{\delta\}_{I}\neq 0$. In fact, the condition $\{\delta\}=0$ implies already in the sub-bins, $\{\delta\}_{I}=\alpha\{\Delta_{N}\}_{I}/N_{0}$, with $\alpha$ a constant. One may understand this linear relation as the leading order expansion of $\{\delta\}_{I}$ in terms of $\{\Delta_{N}\}_{I}/N_{0}$.
Figure 3: (a) Linear correlation between $\{\langle p_{T}\rangle_{I}\}$ and $\{N_{c}\}_{I}$ from sub-bins of the central events from hydrodynamic and HIJING models. (b) Negative linear correlation between $\{\delta\}$ and $\Delta_{N}$ shown in the sub-bin averages of $\{\delta\}_{I}$.
$c_{s}^{2}$ sub-bin slope $\{\delta^{3}\}_{c}=0$$\{\delta^{5}\}_{c}=0$ LEOS
PbPb (Hydro, 5.02TeV) $0.123\pm 0.035$ $0.217\pm 0.032$ $0.216\pm 0.041$ 0.222-0.242
pPb (Hydro, 8.16 TeV) $0.176\pm 0.004$ $0.292\pm 0.013$ $0.287\pm 0.012$ 0.282-0.309
pPb (Hydro, 5.02 TeV) $0.197\pm 0.004$ $0.318\pm 0.011$ $0.313\pm 0.008$ 0.269-0.304
pPb (PYTHIA, 5.02 TeV) $-0.032\pm 0.002$ $1.178\pm 0.006$ $1.352\pm 0.019$ 0.227-0.278
pPb (HIJING, 5.02 TeV) $0.079\pm 0.003$ $1.104\pm 0.019$ $1.171\pm 0.053$ 0.206-0.271
Table 1: A summary of the extracted speed of sound from different methods for various model simulations.
In Fig. 3 (a), using the sub-bin method, the linear correlations between $\{\langle p_{T}\rangle\}_{I}/\langle p_{T}\rangle_{0}$ and $\{N_{\rm c}\}_{I}/N_{0}$ from the different models and systems are shown. Except the results from PYTHIA, where the linear relation has a negative slope, all the model simulations results indicate a positive linear correlation between $\Delta_{p}$ and $\Delta_{N}$. Notably, from the HIJING model, although there is no local thermalization, one still finds a slope that is sizable in magnitude. However, even from hydrodynamic simulations, we find that the values of the slope are substantially small comparing to the physical values of $c_{s}^{2}$ from the LEOS expectations, suggesting a finite and negative contribution from $\{\delta\}_{I}$, which is shown in Fig. 3 (b).
Corresponding to the slope of the sub-bin linearity, and the Gaussianity condition through zero skewness and $\{\delta^{5}\}_{c}=0$, we solve the physical values of the speed of sound . A summary of the obtained results is given in Table. 1. We notice that the extracted values of the speed of sound for all the hydrodynamic model simulations from the Gaussian distribution condition are compatible to the LEOS expectations, within uncertainties. For the non-thermal model simulations, however, the same procedure gives rise to non-physical values of the speed of sound, as they are greater than unity (speed of light $c=1$ as we take the natural units), which violates the causality condition Cherman et al. (2009); Hippert et al. (2025).
Probe of thermalization.– While the exact value of the speed of sound is of great significance to QCD dynamics, we emphasize that the process of extracting the speed of sound also serves as a direct probe of the thermalization of the quark-gluon plasma (QGP), which is also important. The response between mean transverse momentum and particle yield fundamentally differs in a locally thermalized system compared to a non-thermal system. These differences can be effectively distinguished by analyzing the quantum fluctuations of $\delta$ on an event-by-event basis.
Based on our findings, we propose a combined measurement of the speed of sound and the Gaussianity of $\delta$ to characterize local thermalization. In a thermalized system, one expects the extracted speed of sound to have a physical value that falls within a reasonable range Cherman et al. (2009); Hippert et al. (2025), despite uncertainties such as those arising from the determination of the effective temperature, $T_{\rm eff}$, which is partially model-dependent. Simultaneously, the probability distribution $\rho(\delta|c_{s}^{2})$ should exhibit Gaussian behavior, up to corrections of order $1/N_{0}^{a}$. In contrast, if thermalization is not achieved, neither a physical value of the speed of sound nor Gaussianity in $\delta$ should be observed.
Moreover, deviations from thermalization can be quantified using higher-order cumulants of the $\rho(\delta|c_{s}^{2})$ distribution. For example, the standardized kurtosis, defined as $\kappa_{4}=\{\delta^{4}\}/\{\delta^{2}\}^{2}-3$, should vanish in a system with complete local thermalization but remain finite in the presence of non-thermal contributions. From our numerical simulations, we find $\kappa_{4}\gtrsim 2.0$ in both HIJING and PYTHIA models. In contrast, hybrid hydrodynamic modeling yields $\kappa_{4}\approx 0.16$ for p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV and $8.16$ TeV, and $\kappa_{4}\approx-0.06$ for Pb-Pb collisions, indicating small but finite non-thermal contributions.
Summary and discussion. – For a QGP system created in ultracentral nucleus-nucleus collisions, if it is locally thermalized, the response relation between $\Delta_{p}$ and $\Delta_{N}$ is thermodynamic. Accordingly, event-by-event fluctuations, which are primarily quantum in origin, should be independent of the thermodynamic response and follow a Gaussian distribution. Although these fluctuations are correlated with $\Delta_{N}$ and induce corrections to averages, the Gaussianity condition of $\delta$ allows for the statistical extraction of the physical value of the speed of sound, for example, by solving the equation derived from the condition $\{\delta^{3}\}_{c}=0$. We expect the coefficients in this equation to be experimentally measurable, as they correspond to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{N},\Delta_{p})$. Therefore, the physical value of the speed of sound can be precisely determined in experiments, even in systems where fluctuations are substantial, such as in small colliding systems like p-Pb.
The non-Gaussianity of $\delta$ can serve as a robust diagnostic tool for thermalization in QGP systems. In realistic heavy-ion collisions, especially in smaller systems, local thermalization is expected to be partial, with non-thermal contributions becoming more significant in non-central collisions. As a result, we predict a monotonic increase in $\kappa_{4}$ of $\delta$ as a function of centrality. When $\kappa_{4}$ of $\delta$ exceeds unity, the dominant contribution to the response between $\Delta_{p}$ and $\Delta_{N}$ is no longer thermodynamic, indicating that the system should not be considered thermalized. The same argument extends to systems with finite baryon density, where the characterization of thermalization becomes even more crucial for the search for the QCD critical point.
Acknowledgements. – We are grateful to Jean-Yves Ollitrault, Wei Li, and Jiangyong Jia for very helpful discussions. We also thank Xin-Nian Wang for suggesting the HIJING model and Xiangyu Wu for the HIJING code support. Y.-S. M. would like to thank Chen Zhong and Wen-Hao Zhou for their assistance with the computation server. This work is supported in part by the National Natural Science Foundation of China through Grants No. 12375133 and No. 12147101 (L.Y.), No. 12147101, No. 12225502, and No. 12075061 (X.-G. H.), by the Natural Science Foundation of Shanghai through Grant No. 23JC1400200, and by the National Key Research and Development Program of China through Grant No. 2022YFA1604900. The computations are performed at the CFFF platform of Fudan University.
Supplemental Material
The fluid-like fireball created in heavy-ion collisions converts to particles on a hyper-surface $\Sigma$, subject to the freeze-out condition. Through this conversion, the energy and entropy of the expanding fireball are expressed as
[…]
where $\sigma^{\mu}$ denotes the normal vector to the freeze-out hyper-surface, and $T^{\mu\nu}$ and $s^{\mu}$ are the relativistic energy-momentum tensor and entropy flow vector of the fluid, respectively. With respect to the locally equilibrated equation of state (LEOS) that relates energy or entropy density to temperature, the fireball can be effectively reduced to a uniform system characterized by the effective temperature $T_{\rm eff}$ and effective volume $V_{\rm eff}$, as described by the second equation in Eq. (6).
Particles generated from the freeze-out hyper-surface evolve further according to the UrQMD model. The resulting mean transverse momentum $\langle p_{T}\rangle$ of these final particles is expected to exhibit a linear correlation with the effective temperature $T_{\rm eff}$ on an event-by-event basis. For Pb-Pb collisions, which are approximately 2+1 dimensional due to longitudinal boost invariance, we find that $\langle p_{T}\rangle^{(i)}\approx 3T_{\rm eff}^{(i)}$, where $i$ labels the event.
In the p-Pb system, however, longitudinal boost invariance no longer holds, introducing ambiguity in the total energy determined by Eq. (6). The effective temperatures obtained using the $(t,x,y,z)$ coordinate system and the $(\tau,x,y,\eta_{s})$ Milne coordinate system—where $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta_{s}=\tanh^{-1}(z/t)$—are found to differ. Despite this discrepancy, solutions from both coordinate systems exhibit a strong linear correlation with $p_{T}^{(i)}$ on an event-by-event basis, although the correlation coefficients differ. This difference implies that, in a fireball without boost invariance, the mean transverse momentum of the final-state particles may receive contributions from the kinematic energy associated with longitudinal expansion. As there are two sets of linear relations to the effective temperatures, we consider the realistic energy scale of the fireball that is characterized by the mean transverse momentum in the extraction of the speed of sound lies in between, which gives rise to a range in the linear relation between $\langle p_{T}\rangle^{(i)}$ and $T_{\rm eff}^{(i)}$.